# Nonclassical microwave radiation from the parametric dynamical Casimir effect in the reversed-dissipation regime of circuit optomechanics

### v1

@@ VI Nonclassicality of the generated DCE Photons in Phase Space
Beyond the investigation of the Mandel parameter and quadrature squeezing, the generation of nonclassical states emerges as a noteworthy quantum feature of the proposed optomechanical system. In what follows, we examine the potential for generating nonclassical states of Casimir photons enabled by the intrinsic Kerr-type optomechanical nonlinearity inherent in the system operating within the RDR. This capability is fundamentally linked to the presence of the Kerr interaction term $g_{k}\hat{a}^{\dagger 2}\hat{a}^{2}$, which is known to facilitate the generation of Yurke-Stoler states when the system evolves from an initial coherent state over a characteristic time scale $\tau=\pi/2g_{k}$. Notably, coherent states themselves can be efficiently prepared from an initial vacuum state via displacement-type interactions such as $C_{E}(\hat{a}^{\dagger}-\hat{a})$. Since the full system Hamiltonian contains several additional interaction terms beyond the Kerr nonlinearity, it is reasonable to expect that the exact generation of a Yurke-Stoler cat state may be hindered. However, the interplay among these terms can still give rise to approximate nonclassical states for the generated microwave Casimir photons.
The signature of the nonclassical characteristic of the generated states can be revealed by calculating their corresponding Wigner function in the position-momentum phase space which is defined asschleich
@@
Figure 7: Three dimensional plots of the Wigner function, $W(x,p)$, along with its corresponding contour plots, at the time scale $\tau=\pi/2g_{K}$, plotted versus dimensionless canonical quadratures $x$ and $p$. Each 3D surface plot in (a)–(d) has its corresponding contour plot placed directly below it in (e)–(h), respectively. The parameter sets used are: (a,e) $C_{K}=0.05$, $\widetilde{C}_{\epsilon}=0.1$; (b,f) $C_{K}=0.1$, $\widetilde{C}_{\epsilon}=0.1$; (c,g) $C_{K}=10^{-3}$, $\widetilde{C}_{\epsilon}=0.02$; and (d,h) $C_{K}=10^{-3}$, $\widetilde{C}_{\epsilon}=0.03$. Other system parameters are fixed as $\omega_{m}/2\pi=5.33~\text{MHz}$, $\bar{\Delta}_{c}=\omega_{m}$, $C_{E}=10^{-2}$, and $\kappa=0$.
@@
To investigate the influence of system dissipation on the Wigner functions of the generated nonclassical states, and to understand the degree of their nonclassicality, we use the Wigner negativity $\mathcal{W}$, which is defined as kenfack2004
@@
The experimental realization of the setup considered in Fig. 1, has been reported in Ref. Toth2017 . The mechanical resonator contains a parallel-plate capacitor with resonance frequency $\omega_{m}/2\pi=5.33\text{MHz}$ with decay rate of mechanical mode $\gamma_{m}/2\pi=30\text{Hz}$, and effective decay rate $\Gamma_{m}/2\pi=500\text{kHz}$. Also, the total decay rate of primary mode is $\kappa/2\pi=118\text{kHz}$ which is the combination of internal and external decays as $\kappa^{\rm in}/2\pi=76\text{kHz}$ and $\kappa^{\rm ex}/2\pi=42\text{kHz}$, respectively. The laser frequency is approximately $\omega_{L}\approx 1\text{GHz}$.
One of the key challenges in our proposal is determining the value of the optomechanical coupling. By selecting $\bar{\Delta}_{c}=\omega_{m}$, the coupling ratio is expressed as $C_{K}=(g_{0}/\omega_{m})^{2}$, and the values of this parameter we have considered lie in the range $C_{K}\approx 10^{-4}-1$. The vacuum electromechanical coupling strength can be derived from the formula Toth2017 ; thesis-bernier :
@@
Here, $P_{\rm in}$ represents the pump power from the microwave source attenuated by an insertion loss or variable microwave attenuator before reaching the cavity. Note that by tuning and inserting attenuators, one can decrease and adjust our desired values for $P_{\rm in}$. The sideband emission, measured as $P^{\rm meas}_{\rm SB}$, is amplified by an overall gain. To eliminate uncertainties from unknown variables, a weak calibration tone $P_{\rm cal}$ is introduced, following the same path as the pump tone. This calibration tone is tuned near the sideband at $\omega\approx\omega_{m}$, and the corresponding calibration power $P^{\rm meas}_{\rm cal}$ is measured. The initial coupling strength for the given parameters, as reported in Ref. Toth2017 , is $g_{0}^{\rm int}=2\pi\times 106$ Hz. For different parameter sets to attain the desired range of $C_{K}$ values in this paper, the ratio of the coupling strengths must reach $g_{0}/g_{0}^{\rm int}\approx 10^{3}-10^{6}$.
Figure 9 shows the ratio of $g_{0}/g_{0}^{\rm int}$ as a function of temperature and the ratio of pump power $P_{\rm in}/P_{\rm in}^{\rm int}$, where $P_{\rm in}^{\rm int}$ refers to the initial pump power, which corresponds to $g_{0}^{\rm int}$. As shown in the figure, reducing either the pump power or the temperature (relative to their initial conditions) helps to achieve the required coupling strength $g_{0}$ for the desired regime (the region left side of the black dashed line). Notably, these temperatures are experimentally attainable Cattiaux2021 .
@@
The device is mounted on the base plate of a dilution refrigerator, cooled to a base temperature of approximately 10 mK, and further temperature reduction is required to enhance the optomechanical coupling. Additionally, the microwave input lines are heavily attenuated to suppress residual thermal noise, and filter cavities are employed to eliminate unwanted frequency noise from the applied tones Toth2017 .
@@
Also, the laser amplitude is $E_{L}=53\text{kHz}$ which implies the values of the driving laser power have been set $P_{L}\approx-190\text{dBm}$. Moreover, we have set the weak-modulation parameter as $\epsilon\approx 0.01-0.1$ which implies $\epsilon_{L}\approx 10-100\mu\text{Hz}$. Other parameters, although not directly used in calculations, should be specified to establish their experimental range. The intermode coupling strength $\frac{J}{4\pi}|\omega_{\rm aux}-\omega_{c}|=0.57\text{GHz}$, and also the mechanical mode couples to both the primary and auxiliary cavity modes with an initial vacuum electromechanical coupling strength $g_{0}/2\pi=2\times 60\text{Hz}$. Additionally, the internal and external dissipation rates are, respectively, $\kappa_{\rm aux}^{\rm ex}/2\pi=4233$ kHz, and $\kappa_{\rm aux}^{\rm in}/2\pi=245$ kHz which implies the total decay rate of auxiliary being $\kappa_{\rm aux}/2\pi=4487\text{kHz}$, which clearly shows that the critical condition of $\kappa\ll\kappa_{\rm aux}$ for RDR is satisfied. Finally, the resonance frequency of the primary and auxiliary modes are $\omega_{c}/2\pi=4.26\text{GHz}$ and $\omega_{\rm aux}=5.48\text{GHz}$, respectively. It can be noted that there is another manufactured device with different experimental values which is reported in Ref. Toth2017 .
@@
After amplification with a commercial high-electron-mobility transistor (HEMT) amplifier mounted on the $3~K$ plate, the signal can be measured with an electromagnetic spectrum analyzer (ESA) or a vector network analyzer (VNA). Based on the optomechanical sideband cooling by pumping the auxiliary mode corresponding to a mean intracavity photon number $n_{aux}\sim 1.5\times 10^{8}$ on the lower motional sideband, the mechanical oscillator can be prepared as a strongly dissipative-cold reservoir, while still remaining in the weak-coupling regime for the auxiliary mode Toth2017 . Although this order of photons can elevate the mechanical cavity’s temperature and increase thermal photons, this loss can be mitigated Yousefi2023 . Also, a dissipative mechanical reservoir for the primary, high-Q mode is realized since $\Gamma_{m}\gg\kappa$.
@@ VIII Conclusion Remarks
Finally, it should be pointed out that although there are many theoretical DCE proposals as well as some optomechanical proposals for realizing the mechanical DCE, but the present work overcomes the challenges associated with the requirement of the high-frequency MO and with its weak radiation effects.
Acknowledgements.
AMF and HS would thank the ICQT and the CQST, respectively, for their supports. Contributions. AMF proposed and developed the primary idea of DCE in RDR-optomechanics. Also, AMF performed the analytical solution in the short approximation, and then have been checked by HS. MHN provided the method for finding the analog full Hamiltonian of (18). Analytical calculation for finding analog Hamiltonian is performed by HS. All numerical calculations have been done by HS. Experimental discussion is developed by AMF, completed by HS and checked by AY. AY checked the suggested experimental scheme based on the current feasible technologies. RR contributed to discussion related to analog DCE Hamiltonian of the system. The initial manuscript has been written by HS and AMF, and revised by MHN. All authors contributed to prepare the manuscript.
@@ Appendix A Derivation of the Hamiltonian (5)
where $g^{\prime}=g_{0}\bar{a}_{s}$ is the enhanced optomechanical coupling, and $\hat{a}^{\rm aux}_{\rm in}$ and $\hat{b}_{\rm in}$ denote, respectively, the input vacuum noise for the auxiliary mode and the Brownian noise associated with the coupling of the MO to the thermal environment, characterized by the following non-vanishing correlation functions Gardiner2000 .
@@
are, respectively, referred to as the shifted mechanical frequency and effective mechanical damping rate due to the optomechanical interaction Aspelmeyer2014 . Hence, the auxiliary mode with higher damping rate renormalizes both the frequency and the damping rate of the MO.
@@
The maximum amount of optomechanical cooling can be reached in the resolved-sideband regime where $\omega_{m}~\gg~\kappa_{\rm aux}$. In this regime the induced optomechanical frequency and induced optomechanical damping rate are respectively, $\omega_{\rm op}/\omega_{m}\simeq-0.5(g^{\prime}/\omega_{m})^{2}$ and $\Gamma_{\rm op}/\kappa_{\rm aux}\simeq 4(g^{\prime}/\kappa_{\rm aux})^{2}$. Therefore, the shifted mechanical frequency and the effective damping rate in the resolved-sideband regime are, respectively, given by
@@ Appendix B Time evolution operator in the WCR
In order to derive an analytical expression for the system time-evolution operator in the WCR and in the absence of system dissipation, we first transform the Hamiltonian (27) to a frame defined by the unitary transformation $U_{0}={\rm exp}(-i\bar{\Delta}_{c}t\hat{n})$,
@@
where the last term does not influence the dynamics of the system except for a global phase factor $e^{-ig_{K}t/2}$. The first three terms satisfy the su(1,1) Lie algebra with commutation relations $[\hat{L}_{-},\hat{L}_{+}]=2\hat{L}_{z}$ and $[\hat{L}_{z},\hat{L}_{\pm}]=\pm\hat{L}_{\pm}$. For an infinitesimally short interval of time $\delta t$, the time evolution operator is
@@
where $t_{j}=j\delta t$ and $\mathcal{\hat{T}}$ is the time ordering operator. The time dependence of the generators $\hat{L}_{\pm}(t)$ through the exponential terms $e^{\pm 4ig_{K}\hat{n}t}$ makes complicate to find an efficient way to apply the time evolution operator $\hat{U}_{I}(t)$ on an arbitrary initial state. To circumvent this difficulty, we take Roman2015 $e^{\pm 4ig_{K}\expectationvalue{\hat{n}}t}\!\approx 1$, which is justified for times such that $t\ll(4g_{K}\expectationvalue{n})^{-1}$. Under this approximation, the generators $\hat{L}_{\pm}$ appeared in the effective Hamiltonian (63) become time-independent. Now, by applying the Wei-Norman theorem Norman1964 we can express the time evolution operator in the form of a product of exponentials as follows
@@
To derive the analytic expression (31) for the mean number of generated Casimir photons in the WCR and in the absence of cavity dissipation, we need to determine the time-dependence of the operators $\hat{a}(t)$ and $\hat{a}^{\dagger}(t)$. For this purpose, we consider the generalized non-unitary squeezing operators $\hat{S}(\xi,\eta,\zeta)$ defined by Wunsche2003
@@
$[\hat{S}(\xi,\eta,\zeta)]^{\dagger}=\hat{S}(-\xi^{*},-\eta^{*},-\zeta^{*})=[\hat{S}(\xi^{*},\eta^{*},\zeta^{*})]^{-1}.$ (74)

### v2

@@ VI Nonclassicality of the generated DCE Photons in Phase Space
[…]
The signature of the nonclassical characteristic of the generated states can be revealed by calculating their corresponding Wigner function in the position-momentum phase space which is defined as[133]
@@
Figure 7: Three dimensional plots of the Wigner function, $W(x,p)$, along with its corresponding contour plots, at the time scale $\tau=\pi/2g_{K}$, plotted versus dimensionless canonical quadratures $x$ and $p$. Each 3D surface plot in (a)–(d) has its corresponding contour plot placed directly below it in (e)–(h), respectively. The parameter sets used are: (a,e) $C_{K}=0.05$, $\widetilde{C}_{\epsilon}=0.1$; (b,f) $C_{K}=0.1$, $\widetilde{C}_{\epsilon}=0.1$; (c,g) $C_{K}=10^{-3}$, $\widetilde{C}_{\epsilon}=0.02$; and (d,h) $C_{K}=10^{-3}$, $\widetilde{C}_{\epsilon}=0.03$. Other system parameters are fixed as $\omega_{m}/2\pi=5.33\penalty 10000\ \text{MHz}$, $\bar{\Delta}_{c}=\omega_{m}$, $C_{E}=10^{-2}$, and $\kappa=0$.
@@
To investigate the influence of system dissipation on the Wigner functions of the generated nonclassical states, and to understand the degree of their nonclassicality, we use the Wigner negativity $\mathcal{W}$, which is defined as [134]
@@
The experimental realization of the setup considered in Fig. 1, has been reported in Ref. [120]. The mechanical resonator contains a parallel-plate capacitor with resonance frequency $\omega_{m}/2\pi=5.33\text{MHz}$ with decay rate of mechanical mode $\gamma_{m}/2\pi=30\text{Hz}$, and effective decay rate $\Gamma_{m}/2\pi=500\text{kHz}$. Also, the total decay rate of primary mode is $\kappa/2\pi=118\text{kHz}$ which is the combination of internal and external decays as $\kappa^{\rm in}/2\pi=76\text{kHz}$ and $\kappa^{\rm ex}/2\pi=42\text{kHz}$, respectively. The laser frequency is approximately $\omega_{L}\approx 1\text{GHz}$.
One of the key challenges in our proposal is determining the value of the optomechanical coupling. By selecting $\bar{\Delta}_{c}=\omega_{m}$, the coupling ratio is expressed as $C_{K}=(g_{0}/\omega_{m})^{2}$, and the values of this parameter we have considered lie in the range $C_{K}\approx 10^{-4}-1$. The vacuum electromechanical coupling strength can be derived from the formula [120, 135]:
@@
Here, $P_{\rm in}$ represents the pump power from the microwave source attenuated by an insertion loss or variable microwave attenuator before reaching the cavity. Note that by tuning and inserting attenuators, one can decrease and adjust our desired values for $P_{\rm in}$. The sideband emission, measured as $P^{\rm meas}_{\rm SB}$, is amplified by an overall gain. To eliminate uncertainties from unknown variables, a weak calibration tone $P_{\rm cal}$ is introduced, following the same path as the pump tone. This calibration tone is tuned near the sideband at $\omega\approx\omega_{m}$, and the corresponding calibration power $P^{\rm meas}_{\rm cal}$ is measured. The initial coupling strength for the given parameters, as reported in Ref. [120], is $g_{0}^{\rm int}=2\pi\times 106$ Hz.
For different parameter sets to attain the desired range of $C_{K}$ values in this paper, the ratio of the coupling strengths must reach $g_{0}/g_{0}^{\rm int}\approx 10^{3}-10^{6}$. Focusing on the most relevant parameter regime for our results, we consider $g_{0}/g_{0}^{\rm int}\sim 10^{3}\!-\!10^{4}$, which appears more accessible using current state-of-the-art circuit–optomechanical platforms. Importantly, the nonclassical features we predict already emerge in this window: the negative Mandel $Q$ and negative Wigner-function regions (Figs. 4 and 7) occur at $C_{K}=0.05$ ($g_{0}/g_{0}^{\rm int}\approx 10^{4}$, $g_{0}\approx 10^{6}\,\mathrm{Hz}$), while the squeezed states in Fig. 6 appear at $C_{K}=10^{-3}$ ($g_{0}/g_{0}^{\rm int}\approx 10^{3}$, $g_{0}\approx 10^{5}\,\mathrm{Hz}$). These values indicate that the predicted nonclassical behavior is attainable with feasible experimental parameters, which naturally raises the question of how the required ratio $g_{0}/g_{0}^{\rm int}$ can be realized in practice. Firstly, through understanding the system parameters that govern this ratio, and secondly, through practical strategies that directly increase $g_{0}$ in realistic circuit–optomechanical devices.
To address the first aspect, Fig. 9, obtained using Eq. (39), illustrates how the ratio $g_{0}/g_{0}^{\rm int}$ depends on both the temperature and the pump power $P_{\rm in}/P_{\rm in}^{\rm int}$, where $P_{\rm in}^{\rm int}$ is defined by the intrinsic coupling $g_{0}^{\rm int}$. As shown, reducing either the pump power or the temperature helps to reach the required coupling strength $g_{0}$ for accessing the desired regime (region left of the black dashed line). The key parameter controlling $g_{0}/g_{0}^{\rm int}$ is the temperature: ultralow temperatures well below $10\,\mathrm{mK}$, already achievable in modern cryogenic systems, provide favorable conditions to obtain $g_{0}/g_{0}^{\rm int}\sim 10^{3}\!-\!10^{4}$, which is sufficient for observing several of the nonclassical effects discussed in this work [136].
Secondly, as a direct way to increase $g_{0}$, ongoing efforts in mm-wave optomechanical circuits exploit higher cavity frequencies [137, 138, 139]. In addition, experimental groups are exploring high-mechanical-stress superconducting films, such as Niobium and Niobium Titanate, with stresses exceeding 1 GPa. By reducing the capacitor gap below 10 nm, these approaches can achieve up to a 100-fold increase in the intrinsic coupling rate $g_{0}$, providing a complementary path to reach stronger coupling regimes [140]. Together with precise control of temperature, pump power, and device fabrication parameters, these strategies offer practical routes to access stronger coupling regimes and to observe the nonclassical phenomena predicted in this work.
@@
The device is mounted on the base plate of a dilution refrigerator, cooled to a base temperature of approximately 10 mK, and further temperature reduction is required to enhance the optomechanical coupling. Additionally, the microwave input lines are heavily attenuated to suppress residual thermal noise, and filter cavities are employed to eliminate unwanted frequency noise from the applied tones [120].
@@
Also, the laser amplitude is $E_{L}=53\text{kHz}$ which implies the values of the driving laser power have been set $P_{L}\approx-190\text{dBm}$. Moreover, we have set the weak-modulation parameter as $\epsilon\approx 0.01-0.1$ which implies $\epsilon_{L}\approx 10-100\mu\text{Hz}$. Other parameters, although not directly used in calculations, should be specified to establish their experimental range. The intermode coupling strength $\frac{J}{4\pi}|\omega_{\rm aux}-\omega_{c}|=0.57\text{GHz}$, and also the mechanical mode couples to both the primary and auxiliary cavity modes with an initial vacuum electromechanical coupling strength $g_{0}/2\pi=2\times 60\text{Hz}$. Additionally, the internal and external dissipation rates are, respectively, $\kappa_{\rm aux}^{\rm ex}/2\pi=4233$ kHz, and $\kappa_{\rm aux}^{\rm in}/2\pi=245$ kHz which implies the total decay rate of auxiliary being $\kappa_{\rm aux}/2\pi=4487\text{kHz}$, which clearly shows that the critical condition of $\kappa\ll\kappa_{\rm aux}$ for RDR is satisfied. Finally, the resonance frequency of the primary and auxiliary modes are $\omega_{c}/2\pi=4.26\text{GHz}$ and $\omega_{\rm aux}=5.48\text{GHz}$, respectively. It can be noted that there is another manufactured device with different experimental values which is reported in Ref. [120].
@@
After amplification with a commercial high-electron-mobility transistor (HEMT) amplifier mounted on the $3\penalty 10000\ K$ plate, the signal can be measured with an electromagnetic spectrum analyzer (ESA) or a vector network analyzer (VNA). Based on the optomechanical sideband cooling by pumping the auxiliary mode corresponding to a mean intracavity photon number $n_{aux}\sim 1.5\times 10^{8}$ on the lower motional sideband, the mechanical oscillator can be prepared as a strongly dissipative-cold reservoir, while still remaining in the weak-coupling regime for the auxiliary mode [120]. Although this order of photons can elevate the mechanical cavity’s temperature and increase thermal photons, this loss can be mitigated [141]. Also, a dissipative mechanical reservoir for the primary, high-Q mode is realized since $\Gamma_{m}\gg\kappa$.
@@ VIII Conclusion Remarks
For comparison, we note that earlier studies have typically reported either sub-Poissonian photon statistics or Wigner-function negativity, but not both simultaneously within the same physical regime. For example, sub-Poissonian statistics and quadrature squeezing were investigated in vibrating or resonantly oscillating cavities without addressing the Wigner-function negativity [142, 143]. On the other hand, it has been shown [144, 145] that the quantum states of the generated photons in the optical and superconducting analogues of the DCE exhibit Wigner-function negativity, while they obey the (classical) super-Poissonian counting statistics. This comparison highlights one the distinctive aspects of the present work in demonstrating the coexistence of Wigner-function negativity and sub-Poissonian statistics, as nonclassical signatures, within the same physical regime.
Finally, although the effective Hamiltonian studied in this work resembles that of a Kerr nonlinear resonator, its realization in the reversed-dissipation regime of an optomechanical system provides several important advantages. Unlike conventional Kerr platforms, the effective Kerr coefficient in our scheme is tunable in situ through optomechanical parameters, allowing access to regimes of strong nonclassicality—such as simultaneous negativity of the Mandel parameter and the Wigner function—that are not available in standard implementations. This tunability, together with the integrability and versatility of circuit optomechanical devices, makes our platform a promising candidate for generating nonclassical microwave radiation for quantum computing and quantum microwave sensing applications.
Acknowledgments
AMF and HS would thank the ICQT and the CQST, respectively, for their supports. Contributions. AMF proposed and developed the primary idea of DCE in RDR-optomechanics. Also, AMF performed the analytical solution in the short approximation, and then have been checked by HS. MHN provided the method for finding the analog full Hamiltonian of (18). Analytical calculation for finding analog Hamiltonian is performed by HS. All numerical calculations have been done by HS. Experimental discussion is developed by AMF, completed by HS and checked by AY. AY checked the suggested experimental scheme based on the current feasible technologies. RR contributed to discussion related to analog DCE Hamiltonian of the system. The initial manuscript has been written by HS and AMF, and revised by MHN. All authors contributed to prepare the manuscript and revised version based on the referees comments.
@@ Appendix A Derivation of the Hamiltonian (5)
where $g^{\prime}=g_{0}\bar{a}_{s}$ is the enhanced optomechanical coupling, and $\hat{a}^{\rm aux}_{\rm in}$ and $\hat{b}_{\rm in}$ denote, respectively, the input vacuum noise for the auxiliary mode and the Brownian noise associated with the coupling of the MO to the thermal environment, characterized by the following non-vanishing correlation functions [146].
@@
are, respectively, referred to as the shifted mechanical frequency and effective mechanical damping rate due to the optomechanical interaction [48]. Hence, the auxiliary mode with higher damping rate renormalizes both the frequency and the damping rate of the MO.
@@
The maximum amount of optomechanical cooling can be reached in the resolved-sideband regime where $\omega_{m}\penalty 10000\ \gg\penalty 10000\ \kappa_{\rm aux}$. In this regime the induced optomechanical frequency and induced optomechanical damping rate are respectively, $\omega_{\rm op}/\omega_{m}\simeq-0.5(g^{\prime}/\omega_{m})^{2}$ and $\Gamma_{\rm op}/\kappa_{\rm aux}\simeq 4(g^{\prime}/\kappa_{\rm aux})^{2}$. Therefore, the shifted mechanical frequency and the effective damping rate in the resolved-sideband regime are, respectively, given by
@@ Appendix B Time evolution operator in the WCR
[…]
@@
where the last term does not influence the dynamics of the system except for a global phase factor $e^{-ig_{K}t/2}$. The first three terms satisfy the su(1,1) Lie algebra with commutation relations $\hat{L}_{-},\hat{L}_{+}]=2\hat{L}_{z}$ and $\hat{L}_{z},\hat{L}_{\pm}]=\pm\hat{L}_{\pm}$. For an infinitesimally short interval of time $\delta t$, the time evolution operator is
@@
where $t_{j}=j\delta t$ and $\mathcal{\hat{T}}$ is the time ordering operator. The time dependence of the generators $\hat{L}_{\pm}(t)$ through the exponential terms $e^{\pm 4ig_{K}\hat{n}t}$ makes complicate to find an efficient way to apply the time evolution operator $\hat{U}_{I}(t)$ on an arbitrary initial state. To circumvent this difficulty, we take [147] $e^{\pm 4ig_{K}\expectationvalue{\hat{n}}t}\!\approx 1$, which is justified for times such that $t\ll(4g_{K}\expectationvalue{n})^{-1}$. Under this approximation, the generators $\hat{L}_{\pm}$ appeared in the effective Hamiltonian (63) become time-independent. Now, by applying the Wei-Norman theorem [124] we can express the time evolution operator in the form of a product of exponentials as follows
@@
To derive the analytic expression (31) for the mean number of generated Casimir photons in the WCR and in the absence of cavity dissipation, we need to determine the time-dependence of the operators $\hat{a}(t)$ and $\hat{a}^{\dagger}(t)$. For this purpose, we consider the generalized non-unitary squeezing operators $\hat{S}(\xi,\eta,\zeta)$ defined by [148]
@@
$\hat{S}(\xi,\eta,\zeta)]^{\dagger}=\hat{S}(-\xi^{*},-\eta^{*},-\zeta^{*})=[\hat{S}(\xi^{*},\eta^{*},\zeta^{*})]^{-1}.$ (74)